# Superposition in Transformers: A Novel Way of Building Mixture of Experts

### v1

@@ 1 Introduction
Large language models (LLMs) such as GPT-3[1] and GPT-4[2] have shown remarkable performance on various language tasks. However, adapting these models to new tasks or domains often leads to catastrophic forgetting, where newly learned information overwrites older knowledge [3]. This challenge becomes critical in continual learning or domain adaptation settings, where preserving performance on original tasks is essential [4].
Mixture of Experts (MoE) [5] models mitigate catastrophic forgetting by delegating different distributions or tasks to distinct experts. However, MoE solutions typically require large parameter expansions and introduce complex gating. In contrast, we propose a more parameter-efficient method: we blend the hidden states of two models—a base model and a fine-tuned model—using trainable blending coefficients derived from B-splines. We then reconstruct the original state from the blended weights via the autoencoders at inference time, allowing each model’s states to co-exist in a superposed fashion with minimal interference.
1.1 Contributions
@@
1.2 Motivation and Positioning
Most existing mixture-of-experts methods expand parameter count significantly by introducing distinct expert modules and gating networks [6, 7]. Parameter-efficient fine-tuning approaches such as Adapters [8] or LoRA [9] often rely on modifying or appending new weights to the base model. In contrast, our method produces a single merged set of parameters—via B-spline blending—while training autoencoders to reconstruct the original hidden states on demand. During training, both the base and fine-tuned model parameters are accessed in a frozen manner to generate hidden states, but in the end, only the new blended model parameters (plus the autoencoders) remain. This design merges the capabilities of two experts into one compact parameter space and mitigates catastrophic forgetting by preserving each model’s knowledge in a superposed, reconstructable form.
@@ 2 Background and Related Work
2.1 Mixture of Experts
Mixture of Experts (MoE) dynamically allocates different experts to different inputs, improving capacity and specialization. While effective, MoE methods often increase parameter count and rely on gating networks [7]. Our method similarly aims to leverage multiple “expert” models but does so via a shared parameter space and lightweight autoencoder modules.
2.2 Parameter-Efficient Transfer Learning
Recent techniques like Adapters [8] and LoRA [9] reduce the cost of fine-tuning large models by injecting small trainable components. We differ in that we freeze both the base and fine-tuned models; instead of adding new layers directly on top of them, we blend their hidden states and reconstruct them adaptively. This approach can be viewed as orthogonal to Adapters or LoRA, potentially combining well with these methods.
2.3 Superposition and Polysemantic Neurons
In neural networks, superposition refers to efficiently encoding multiple features or tasks within the same parameter space [10]. Large LLMs often exhibit polysemantic neurons—units that respond to multiple unrelated concepts [11]. Our approach further promotes polysemanticity by forcing the network to encode the states of two distinct models in overlapping neurons, constrained by autoencoder bottlenecks.
2.4 Neural Network Compression and Model Merging
Prior methods for merging or compressing models [12, 13] generally aim to reduce memory overhead. By contrast, our framework focuses on preserving both original models in a single parameter space without overwriting. The added overhead is small (primarily the autoencoders and the blending coefficients) compared to training or storing two separate models.
@@ 3.1 Overview
3.2 Blending Hidden States Using B-Splines
@@ 3.2.2 Formulation
  • $\{c_{i}\}$ are trainable control points for a B-spline of degree $k$.
  • $b_{l}$ is a layer-specific bias term.
  • The B-spline basis functions $B_{i,k}(\cdot)$ ensure smoothly varying $\alpha(l)$.
We freeze the weights of both $M_{\text{base}}$ and $M_{\text{fine}}$, focusing on learning only $\{c_{i}\}$, $\{b_{l}\}$, and the autoencoders.
@@ 3.3 Merged Model Architecture
3.3.1 Merging Post-Training
@@ 3.4.1 Architecture
At each selected layer $l$, an autoencoder compresses the blended hidden state $\mathbf{h}_{l}$ into a bottleneck and reconstructs it:
$\mathbf{z}_{l}=\mathrm{Encoder}(\mathbf{h}_{l}),\quad\hat{\mathbf{h}}_{l}=\mathrm{Decoder}(\mathbf{z}_{l}).$
@@ 3.4.4 Extending to a 2D-alpha Model (Optional)
One can extend $\alpha(l)$ to a vector for per-dimension blending. In this scenario, local features (extracted by convolution layers) and global features (captured by a low-rank adapter) are combined in the autoencoder. While more expressive, this 2D-alpha approach also increases complexity and may demand careful tuning to avoid overfitting.
4 Training Procedure
4.1 Objectives
We optimize two core objectives:
  • $\mathcal{L}_{\mathrm{LM}}$: Standard language modeling loss, preserving fluency and coherence.
  • $\mathcal{L}_{\mathrm{Recon}}$: Reconstruction error between $\hat{\mathbf{h}}_{l}$ and the target hidden state $\mathbf{h}_{l}^{\text{base}}$ or $\mathbf{h}_{l}^{\text{fine}}$.
An additional alpha regularization term $\mathcal{L}_{\mathrm{AlphaReg}}$ may be applied to keep $\alpha$ smooth or centered.
4.2 Loss Functions
@@
$\mathcal{L}\;=\;\lambda_{\mathrm{LM}}\,\mathcal{L}_{\mathrm{LM}}\;+\;\lambda_{\mathrm{Recon}}\,\mathcal{L}_{\mathrm{Recon}}\;+\;\lambda_{\mathrm{Alpha}}\,\mathcal{L}_{\mathrm{AlphaReg}},$
where $\lambda_{\mathrm{LM}}$, $\lambda_{\mathrm{Recon}}$, and $\lambda_{\mathrm{Alpha}}$ balance the relative importance of each component. $\mathcal{L}_{\mathrm{Recon}}$ is typically Mean Squared Error (MSE) or L2 distance over the hidden states.
4.3 Optimization
@@ 5 Experiments and Results
In this section, we evaluate the proposed merged model, by integrating a base GPT-2 model and a fine-tuned GPT-2 model trained on French text. We focus on demonstrating how the autoencoders enable the superposition of transformer blocks from different fine-tunes, allowing the merged model to effectively combine representations from both models. Our analysis includes performance metrics, hidden state reconstruction and the emergence of polysemantic neurons.
@@ 5.1 Experimental Setup
We compare the following models:
  • Base Model: GPT-2 trained on English data.
  • Fine-Tuned Model: GPT-2 fine-tuned on a French corpus.
  • Merged Model: Combines the base and fine-tuned models using layer-wise blending with learned $\alpha$ values and incorporates autoencoders to enable superposition.
Datasets
English and French evaluation datasets:
  • English Dataset (DS1): A subset of the GPT-2 training data (6,000 samples repeated with range(3) in the 1D experiment and repeated with range(2) in the 2D experiment).
  • French Dataset (DS2): 18,000 French Wikipedia articles for the 1D model and 12,000 French Wikipedia articles for the 2D model.
10% of the combined samples are reserved for validation.
In the case of the 1D-alpha GPT2 model the number of parameters per autoencoder is:
@@
The bottleneck size used in the 1D-alpha model experiment is 576, resulting in approximately 1.7 million parameters per autoencoder.
@@ 5.3 Results of the 2D model
5.3.1 Visualization of Hidden States in Layer 4
@@ 5.3.2 Neuron Utilization and Polysemantic Neurons
The merged model maintains efficient neuron utilization and shows an increase in polysemantic neurons, as shown in Table 6:
@@
The 2D-alpha model demonstrates enhanced neuron utilization. The increase in sparsity is minimal and potentially arbitrary due to the model’s small size.
@@ 5.3.3 Comparative Analysis of Neuron Diversity and Activation
Note: In the 2D-alpha model, the autoencoders were used starting from layer 4 until layer 10. The graphs below clearly illustrate the contrast between layer 3 and the subsequent layers 4, 5, and 6.
@@
  • The reduction in diversity metrics shows that neurons encode overlapping or generalized features. While layers without (e.g. layer 3) autoencoders retain higher diversity due to the presence of specialized neurons, the autoencoder forces neurons to unify and encode features relevant across multiple tasks or contexts. This suggests a shift from task-specificity to multi-purpose functionality—a hallmark of polysemantic neurons.
@@ 7 Conclusion
Appendix A Conventions
In this paper, we refer to the original, unmodified model as the base model, denoted by $M_{\text{base}}$ and to the fine-tuned version, referred to as $M_{\text{fine}}$. We have implemented two distinct architectures: the first serves as a simplified implementation (in terms of number of parameters) to demonstrate the aforementioned properties.This model refered to as the ”1D-alpha model” due to its use of scalar alpha values for layer-wise adjustments. The second delves deeper into exploring the impact of the autoencoder on the polysemantic nature of neurons within a transformer block and is refered to as the ”2D-alpha model” due to its utilization of vector-based alpha values for each layer. Collectively, any form of these models will be referred to as the merged model $M_{\text{merged}}$.
$L$ denotes the number of transformer blocks within the model, and $D$ is the hidden size of each transformer block.
Appendix B Training Algorithms
B.1 Loss Functions
@@
B.2 1D-alpha Model Training algorithm
@@
B.3 2D-alpha Model Training algorithm
@@
Appendix C Additional results of the 1D-alpha Model experiment
Tables 1 and 2 compare the perplexity of the models using French and English inputs :
Model Language Perplexity
Base Model English 28.88
French 132.02
Fine-Tuned Model English 54.42
French 29.57
Merged Model English 43.14
French 33.20
Table 2: Perplexity of the models on the evaluation datasets
The perplexity results show distinct performance characteristics for the models. For French inputs, the fine-tuned model has a significantly lower perplexity (29.57) compared to the base model (132.02), indicating its suitability for French. Conversely, the base model performs best on English inputs with a lower perplexity (28.88), suggesting it was primarily optimized for English. The merged model demonstrates comparable perplexity to the base model in English (43.14 vs. 28.88) and is close to the fine-tuned model in French (33.20 vs. 29.57), indicating its effectiveness in bridging the performance across both languages.
C.1 Visualization of Hidden States in Layer 4
@@
The t-SNE projection in Figure 2 provides a visual representation of how the merged model’s layer 4 hidden states are reconstructed based on the input data. Specifically:
@@
C.2 Evaluation of Hidden State Trajectories

### v2

@@ 1 Introduction
Large language models (LLMs) such as GPT-3 [1] and GPT-4 [2] have shown remarkable performance on various language tasks. However, adapting these models to new tasks or domains often leads to catastrophic forgetting, where newly learned information overwrites older knowledge [3]. This challenge becomes critical in continual learning or domain adaptation settings, where preserving performance on original tasks is essential [5, 6].
Most existing Mixture of Experts (MoE) [7] methods expand parameter count significantly by introducing distinct expert modules and gating networks [8, 9]. Parameter-efficient fine-tuning approaches such as Adapters [10] or LoRA [11] often rely on modifying or appending new weights to the base model [12, 13, 14]. In contrast, our method produces a single merged set of parameters—via B-spline blending—while training autoencoders to reconstruct the original hidden states on demand. During training, both the base and fine-tuned model parameters are accessed in a frozen manner to generate hidden states, but in the end, only the new blended model parameters (plus the autoencoders) remain. This design merges the capabilities of two experts into one compact parameter space and mitigates catastrophic forgetting by preserving each model’s knowledge in a superposed, reconstructable form.
The contributions of this paper are as follows:
@@ 2 Background and Related Work
Our work lies at the intersection of several key research areas in neural networks, including Mixture of Experts (MoE) models, parameter-efficient transfer learing, and neural network compression. Additionally, we draw from studies on superposition and polysemantic neurons in language models.
Mixture of Experts. MoE models dynamically allocates different experts to different inputs, improving capacity and specialization. While effective, MoE methods often increase parameter count and rely on gating networks [9]. Our method similarly aims to leverage multiple “expert” models but does so via a shared parameter space and lightweight autoencoder modules.
Parameter-Efficient Transfer Learning . Recent techniques like Adapters [10] and LoRA [11] reduce the cost of fine-tuning large models by injecting small trainable components. We differ in that we freeze both the base and fine-tuned models; instead of adding new layers directly on top of them, we blend their hidden states and reconstruct them adaptively. This approach can be viewed as orthogonal to Adapters or LoRA, potentially combining well with these methods.
Superposition and Polysemantic Neurons . In neural networks, superposition refers to efficiently encoding multiple features or tasks within the same parameter space [15]. Large LLMs often exhibit polysemantic neurons—units that respond to multiple unrelated concepts [16]. Recent research into feature interpretability has shown that sparse autoencoders (SAEs) can transform model activations into more interpretable feature spaces, helping to disentangle polysemantic neurons into monosemantic units  [18, 4, 17]. By leveraging these learned features, researchers have identified significant similarities across latent spaces in different LLMs, suggesting a shared representational structure  [18]. Our approach promotes polysemanticity by forcing the network to encode the states of two distinct models in overlapping neurons, constrained by autoencoder bottlenecks.
Neural Network Compression and Model Merging . Prior methods for merging or compressing models [19, 20] generally aim to reduce memory overhead. By contrast, our framework focuses on preserving both original models in a single parameter space without overwriting. The added overhead is small (primarily the autoencoders and the blending coefficients) compared to training or storing two separate models.
Preventing Catastrophic Forgetting . Preventing catastrophic forgetting, where models lose previously learned knowledge when fine-tuned on new tasks, has been a critical area of research in transfer learning [6]. Our method addresses this by freezing the base and fine-tuned models, thus preserving their original knowledge. The B-spline blending coefficients and autoencoders provide a flexible mechanism for combining knowledge from both models without overwriting their respective features. This approach shares similarities with knowledge distillation techniques but focuses on preserving knowledge across layers rather than compressing it into a smaller model.
@@ 3.1 Overview
3.2 Blending Model Weights Using B-Splines
@@ 3.2.2 Formulation
where $\{c_{i}\}$ are trainable control points for a B-spline of degree $k$ and $b_{l}$ is a layer-specific bias term.
The B-spline basis functions $B_{i,k}(\cdot)$ ensure smoothly varying $\alpha(l)$. We freeze the weights of both $M_{\text{base}}$ and $M_{\text{fine}}$, focusing on learning only $\{c_{i}\}$, $\{b_{l}\}$, and the autoencoders.
@@ 3.3 Merged Model Architecture
3.3.1 Merging Models Post-Training
@@ 3.4.1 Architecture
At each selected layer $l$, an autoencoder compresses the blended hidden state $\mathbf{h}_{l}$ into a bottleneck and reconstructs it: $\mathbf{z}_{l}=\mathrm{Encoder}(\mathbf{h}_{l})$ and $\hat{\mathbf{h}}_{l}=\mathrm{Decoder}(\mathbf{z}_{l})$.
@@ 3.4.4 Extending to a 2D-alpha Model (Optional)
One can extend $\alpha(l)$ to a vector for per-dimension blending. In this scenario, local features (extracted by convolution layers) and global features (captured by a low-rank adapter) are combined in the autoencoder. The 2D-alpha approach delves deeper into exploring the impact of the autoencoder on the polysemantic nature of neurons within a transformer block. While more expressive, it also increases complexity and may demand careful tuning to avoid overfitting.
[…]
Figure 1: Overview of a GPT-2 Merged Model Architecture.
3.5 Training Procedure
3.5.1 Objectives
We optimize two core objectives: a standard language modeling loss $\mathcal{L}_{\mathrm{LM}}$, preserving fluency and coherence and the reconstruction error between $\hat{\mathbf{h}}_{l}$ and the target hidden state $\mathbf{h}_{l}^{\text{base}}$ or $\mathbf{h}_{l}^{\text{fine}}$. An additional alpha regularization term $\mathcal{L}_{\mathrm{AlphaReg}}$ may be applied to keep $\alpha$ smooth or centered.
3.5.2 Loss Functions
@@
$\mathcal{L}\;=\;\lambda_{\mathrm{LM}}\,\mathcal{L}_{\mathrm{LM}}\;+\;\lambda_{\mathrm{Recon}}\,\mathcal{L}_{\mathrm{Recon}}\;+\;\lambda_{\mathrm{Alpha}}\,\mathcal{L}_{\mathrm{AlphaReg}}$
where $\lambda_{\mathrm{LM}}$, $\lambda_{\mathrm{Recon}}$, and $\lambda_{\mathrm{Alpha}}$ balance the relative importance of each component. $\mathcal{L}_{\mathrm{Recon}}$ is typically Mean Squared Error (MSE) or L2 distance over the hidden states. $\mathcal{L}_{\mathrm{AlphaReg}}$ encourages the control points of the B-spline-based blending coefficients to adhere to desirable properties. It ensures smooth and interpretable blending, avoiding overfitting to noisy representations.
3.5.3 Optimization
@@ 4 Experiments and Results
In this section, we evaluate the proposed merged model, by integrating a base GPT-2 [21] model and a fine-tuned GPT-2 model trained on French text. We focus on demonstrating how the autoencoders enable the superposition of transformer blocks from different fine-tunes, allowing the merged model to effectively combine representations from both models. Our analysis includes performance metrics, hidden state reconstruction and the emergence of polysemantic neurons.
@@ 4.1 Experimental Setup
We compare a Base Model (GPT-2 trained on English data), a Fine-Tuned Model (GPT-2 fine-tuned on a French corpus) and the Merged Model which combines the base and fine-tuned models using layer-wise blending with learned $\alpha$ values and incorporates autoencoders to enable superposition. For evaluation, we used two datasets, an english dataset that contains a subset of the GPT-2 training data (6,000 samples repeated with range(3) in the 1D experiment and repeated with range(2) in the 2D experiment) and a french dataset (18,000 French Wikipedia articles for the 1D model and 12,000 French Wikipedia articles for the 2D model). 10% of the combined samples are reserved for validation.
@@ 4.2 Results of the 1D model
4.2.1 Perplexity per language
Tables 1 compares the perplexity of the models using French and English inputs : The perplexity results show distinct performance characteristics for the models. For French inputs, the fine-tuned model has a significantly lower perplexity (29.57) compared to the base model (132.02), indicating its suitability for French. Conversely, the base model performs best on English inputs with a lower perplexity (28.88), suggesting it was primarily optimized for English.
The merged model demonstrates closer perplexity to the base model in English (43.14 vs. 28.88) and is comparable to the fine-tuned model in French (33.20 vs. 29.57), indicating its effectiveness in bridging the performance across both languages.
Model Language Perplexity
Base Model English 28.88
French 132.02
Fine-Tuned Model English 54.42
French 29.57
Merged Model English 43.14
French 33.20
Table 1: Perplexity of the models on the evaluation datasets.
The merged model demonstrates closer perplexity to the base model in English (43.14 vs. 28.88) and is comparable to the fine-tuned model in French (33.20 vs. 29.57), indicating its effectiveness in bridging the performance across both languages.
@@ 4.3 Results of the 2D model
4.3.1 Hidden States in Layer 4
@@ 4.3.2 Neuron Utilization and Polysemantic Neurons
The 2D-alpha model demonstrates enhanced neuron utilization (table 2). The increase in sparsity is minimal and potentially arbitrary due to the model’s small size.
@@ 4.3.3 Comparative Analysis of Neuron Diversity and Activation
Note: In the 2D-alpha model, the autoencoders were used starting from layer 4 until layer 10. Figures 4 and 5 clearly illustrate the contrast between layer 3 and the subsequent layers 4, 5, and 6.
The reduction in diversity metrics (figure 4) shows that neurons encode overlapping or generalized features. While layers without (e.g. layer 3) autoencoders retain higher diversity due to the presence of specialized neurons, the autoencoder forces neurons to unify and encode features relevant across multiple tasks or contexts. This suggests a shift from task-specificity to multi-purpose functionality—a hallmark of polysemantic neurons.
@@
5 Limitations
While our method demonstrates promising results, several limitations should be noted. Currently, this approach might does not support dynamic state switching within a single input. Achieving this may require additional mechanisms, such as token-level state signaling, which remains unexplored. Additionally, our experiments focus solely on merging two models—one base model and one fine-tuned model. Extending this approach to handle more than two models, such as combining multiple experts or integrating diverse knowledge domains, may pose scalability challenges and require modifications to the blending and reconstruction processes.
@@ 7 Conclusion
Appendix
Appendix A Training Algorithms
A.1 Loss Functions
@@
A.2 1D-alpha Model Training algorithm
@@
A.3 2D-alpha Model Training algorithm
@@
Appendix B Additional results of the 1D-alpha Model experiment
B.1 Visualization of Hidden States in Layer 4
@@
The t-SNE projection in Figures 6 and 7 provide a visual representation of how the merged model’s layer 4 hidden states are reconstructed based on the input data. Specifically:
@@
B.2 Evaluation of Hidden State Trajectories